OpenStudy (anonymous):

Find the equation of the line in slope intercept form that passes through the coordinates (-1,2) and (7,6).

OpenStudy (anonymous):

What do I do?

OpenStudy (geometry_hater):

Slope intercept.......Y= Mx + B So we dont have a y intercept or a Slope so we need to find both of them First finding the slope....Simply do the formula Y2 - Y1 over X2-X1 and there's your slope.

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

first try to draw it on a graph to visualize it...

OpenStudy (anonymous):

I got 1/2... FOR THE UM...slope

OpenStudy (anonymous):

the slope is positive so 1/2 you are right...

OpenStudy (anonymous):

But then what do I do next?

OpenStudy (anonymous):

now plug in it to find B...use any point ...

OpenStudy (anonymous):

you can use any of the two set points...(-1,2) or (7,6) ....

OpenStudy (anonymous):

so... 2=1/2(-1) + b?

OpenStudy (anonymous):

Got b= 5/2...

OpenStudy (anonymous):

yes...now a little bit of algebra to find B which is the Y intersection..

OpenStudy (anonymous):

x.x...

OpenStudy (anonymous):

So...substitute x w/ 0 to find y...

OpenStudy (anonymous):

not really you already got B....just put it in slope intercept form....y= m(1/2)X+b(5/2)

OpenStudy (anonymous):

So that's it?

OpenStudy (anonymous):

yup.....Y=Mx+B...we found M and B ...so put it into slope intercept form...

OpenStudy (anonymous):

okay, thanks.

OpenStudy (anonymous):

Youre welcome